OpenStudy (anonymous):

solve the equation by using logarithm 7^x-1= 5^2x+1 leave answer with log please show all work

OpenStudy (annon):

(x-1)log7=(2x+1)log5

OpenStudy (annon):

divid a log from either side and solve for x

OpenStudy (anonymous):

apply log on both sides \[\log7^{x-1}=\log(5^{2x}+1)\] \[\log7^{x-1}=\log5^{2x} * log1\] \[\ (x-1)log7=0\] x =1

OpenStudy (anonymous):

wait, is the 2x+1 in the exponent of 5 or is it just 2x?

OpenStudy (anonymous):

2x+1 is the exponents of 5

OpenStudy (anonymous):

oh, then do what annon said.

OpenStudy (anonymous):

can u do it?

OpenStudy (anonymous):

yes, but I prefer you do it.

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

it is not the case that \[log(5^{2x}+1)=log(5^{2x})\times log(1)\]

OpenStudy (anonymous):

is u studying? dhatraditya

OpenStudy (anonymous):

that would make everything 0!

OpenStudy (anonymous):

satellite, I had the same thought, but apparently it is \[5^{2x+1}\]

OpenStudy (anonymous):

doesn't matter it is never the case

OpenStudy (anonymous):

no i'm not studying. Im here to help

OpenStudy (anonymous):

\[7^{x-1}=5^{2x+1}\] yes?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

or (x-1) log 7 = (2x+1) log 5

OpenStudy (anonymous):

or 1.2(x-1) = 2x+1 solve for x

OpenStudy (anonymous):

\[ln(7^{x-1})=ln(5^{2x+1})\] \[(x-1)ln(7)=(2x+1)ln(5)\]

OpenStudy (anonymous):

\[ln(7)x-ln(7)=2ln(5)x+ln(5)\]

OpenStudy (anonymous):

I don't know anything about log. I read in 9th grade. and these problems are from my college now class

OpenStudy (anonymous):

\[ln(7)x-2ln(5)x=ln(5)+ln(7)\]

OpenStudy (anonymous):

\[x(ln(7)-2ln(5))=ln(5)+ln(7)\]

OpenStudy (anonymous):

and finally \[x=\frac{ln(5)+ln(7)}{ln(7)-2ln(5)}\]

OpenStudy (anonymous):

only log step was to get the variable out of the exponent. all else was algebra

OpenStudy (anonymous):

you can write the answer in different ways using propertties of the logs

OpenStudy (anonymous):

for example \[ln(5)+ln(7)=ln(35)\] but i would just leave it

OpenStudy (anonymous):

halimabegum wrote "I don't know anything about log. I read in 9th grade. and these problems are from my college now class" Thats hardly an excuse.

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

problems like these are a pain where you have different bases. you can see all the algebra involved